# Knowledge Graph Large Language Model (KG-LLM) for Link Prediction

### v5

@@ Abstract
The task of predicting multiple links within knowledge graphs (KGs) stands as a challenge in the field of knowledge graph analysis, a challenge increasingly resolvable due to advancements in natural language processing (NLP) and KG embedding techniques. This paper introduces a novel methodology, the Knowledge Graph Large Language Model Framework (KG-LLM), which leverages pivotal NLP paradigms. We use our method to convert structured knowledge graph data into natural language, and then use these natural language prompts to finetune large language models (LLMs) to enhance multi-hop link prediction in KGs. By converting the KG to natural language prompts, our framework is designed to discern and learn the latent representations of entities and their interrelations. To show the efficacy of the KG-LLM Framework, we fine-tune three leading LLMs within this framework, employing both non-ICL and ICL tasks for a comprehensive evaluation. Further, we explore the framework’s potential to provide LLMs with zero-shot capabilities for handling previously unseen prompts. Our experimental findings discover that the performance of our approach can significantly boost the models’ generalization capacity, thereby ensuring more precise predictions in unfamiliar scenarios.
@@ 1 Introduction
In the realm of data representation and organization, knowledge graphs (KGs) have emerged as a structured and effective methodology, attracting substantial interest in recent years. Despite their utility, gaining valuable insights from KGs, especially when navigating the multi-hop entity relationships, poses a considerable challenge. Although two-node link prediction has yielded promising results, the prediction of multi-hop links remains a difficult task. This difficulty stems from the necessity for models to reason about intermediate entities and their relationships. Further challenge is the issue of debugging KG model predictions, particularly in the context of discriminative prediction, where the model’s lack of explanatory reasoning steps, which obscures the origins of errors, diminishing accuracy and performance. Consequently, the development of models capable of generatively and precisely predicting multi-hop links in KGs is a critical challenge.
Figure 1: One example of KG-LLM framework can be visually demonstrated with a simple task consisting of two stages. In the first stage, we apply the graph preprocessing function to a randomly selected path in the knowledge graph, enabling the path to be represented as a chain-of-thought prompt. The second stage use instruction finetuning to finetune the LLM to predict whether the first node is connected to the last node in the sub-KG.
Historically, approaches to solving tasks related to KGs can trace their origins from embedding-based methods to more recent advancements with LLMs [25]. Initially, embedding-based methods were pivotal, utilizing techniques to represent both entities and relations in a KG as low-dimensional vectors to address the link prediction task by preserving the structural and semantic integrity of the graph [2], [27], [5], [9]. As the field progressed, the integration of LLMs began to offer new paradigms, leveraging vast amounts of data and sophisticated architectures to further enhance prediction capabilities and semantic understanding in KGs [1], [35], [36], [34], [19]. This transition shows a significant improvement from using purely mathematical representations to more context-aware methodologies that better understand the knowledge representation.
@@
To bridge these gaps, we propose the Knowledge Graph Large Language Model Framework (KG-LLM Framework), a novel approach to multi-hop link prediction. As illustrated in Figure 1, envision a KG wherein nodes are interconnected through specific relations. Initially, our framework takes input of the original KG dataset. After preprocessing, a path in the KG is transformed into a chain-of-thought (CoT) prompt [30], enclosing through a series of relational statements, which can be represented as {Node 1 (has relation_$x$ with) Node 2, Node 2 (has relation_$y$ with) Node 3, etc.}. The complexity of a KG’s CoT prompt, determined by its path length and the number of nodes and connections, directly correlates with the intricacy of the multi-hop problem it represents. Through instruction fine-tuning (IFT) [28] of three Large Language Models (LLMs)—Flan-T5 [28], LlaMa2 [21], and Gemma [20]—our framework is ready to enhance multi-hop link prediction performance during the testing phase. Moreover, by integrating in-context learning (ICL) [32], the model not only improves but also acquires the capacity to tackle unseen prompts, showcasing our method’s innovativeness in addressing multi-hop link prediction challenges.
Our study presents the KG-LLM framework as a pioneering approach in the realm of multi-hop link prediction. Our key contributions are:
@@
  • Our analysis on real-world datasets confirms that our framework improves generative multi-hop link prediction in KGs, underscoring the benefits of incorporating CoT and instruction fine-tuning during training.
@@ 2 Related Work
Recent development of large-scale language models, such as BERT [4], GPT [14], LLaMA [21], Gemini [20], T5 [15], and Flan-T5 [28] has significantly advanced NLP in many ways. These models are trained on vast amounts of text data and have demonstrated impressive performance on a wide range of NLP tasks. The text-to-text training approach makes LLMs particularly suitable for our generative multi-hop link prediction task, where the model needs to generate the connection between multiple entities based on the input text. Recent studies, concurrent work such as GraphEdit [6], MuseGraph [18] and InstructGraph [24], have shown that natural language is effective for representing structural data for LLMs. Besides, training on large-scale data makes it possible for LLMs to generalize to unseen tasks or prompts that were not part of its training data [29]. In our study, we will utilize an unseen prompt “What is the relation between Node_id and Node_id?” instead of the original prompt “Is Node_id connected with Node_id?” to evaluate the model’s ability to generalize to unseen prompts.
Another advantage of LLM-based generative modeling is the Chain-of-Thought (CoT) reasoning ability [31], which provides the flexibility of modifying the instruction, options, and exemplars to allow structured generation and prediction. The Chain-of-Thought reasoning process can be naturally integrated with knowledge graphs by translating a reasoning path on a knowledge graph into a natural language style. This flexibility allows us to easily test the model’s ability to follow instructions and make decisions based on the provided information.
Similarly, In-Context Learning (ICL) [3] helps large language models learn from demonstrative examples in the prompt to generate correct answers for the given question. This can also be naturally integrated with knowledge graphs by considering multi-path reasoning in KG. As a result, CoT plus ICL together enable very flexible KG reasoning through natural language, where CoT enables path-based reasoning over KG and ICL further extends the power to multiple paths.
Figure 2: Models processed through the ablation framework will be trained using the KG-LLM ablation knowledge prompt dataset (left), whereas models processed via the KG-LLM framework will be trained on the KG-LLM knowledge prompt (right).
@@ 3.1.1 Multi-hop Link Prediction
The task of multi-hop link prediction extends beyond simple edge prediction by aiming to identify missing connections over multiple relational steps within a knowledge graph. Specifically, given a sequence of observed triples that form a connected path $P_{obs}={(e_{1},r_{1},e_{2}),(e_{2},r_{2},e_{3}),\ldots,(e_{n-1},r_{n-1},e_{n})}\subseteq L$, where each triple $(e_{i},r_{i},e_{i+1})$ denotes an observed relation $r_{i}$ between entities $e_{i}$ and $e_{i+1}$. The objective is to predict whether a missing link $l_{miss}=(e_{1},?,e_{n})$ exists by answering True or False [16], [23].
@@ 3.1.2 Multi-hop Relation Prediction
The task of multi-hop relation prediction closely aligns the concept of multi-hop link prediction, with a critical distinction in the question and output. Rather than determining the existence of a missing link $l_{miss}$, this task delves into specifying the relationship. This change shifts the focus from a binary existence query to identifying the specific relationship that binds the entities [12].
@@ 3.2 Knowledge Prompt
The knowledge prompt refers to a specialized prompt designed for KGs that enables the conversion of a given sequence of observed triples $P_{obs}$ into natural language, thereby facilitating the model learn better. By leveraging the knowledge prompt in the training process, the model is able to more effectively understand the underlying relationships and patterns present within KGs, thus improving overall performance in multi-hop link prediction tasks. In this paper, we define the two types of knowledge prompts, KG-LLM knowledge prompt and KG-LLM (ablation) knowledge prompt:
Multi-hop Link Prediction KG-LLM Knowledge Prompt:
###Instruction:
[…]
Answer the question by reasoning step-by-step. Choose from the given options:
1. Yes
2. No
###Input:
[…]
Multi-hop Relation Prediction KG-LLM Knowledge Prompt:
###Instruction:
[…]
Answer the question by reasoning step-by-step. Choose from the given options:
[…]
###Input:
[…]
Multi-hop Link Prediction KG-LLM(ablation) Knowledge Prompt:
###Input:
[…]
Multi-hop Relation Prediction KG-LLM(ablation) Knowledge Prompt:
###Input:
[…]
The two types of prompts demonstrate distinct approaches to enhancing model performance in multi-hop prediction tasks. KG-LLM knowledge prompt adopts a structured format that includes instructions and inputs. This approach involves breaking down complex and lengthy inputs into manageable, concise processes. By organizing the information systematically, the likelihood of the model misinterpreting the queries is significantly reduced. Such a methodical strategy ensures the model stays highly aware of the Chain-of-Thought (CoT) process throughout its training, thereby supporting its ability to interpret and navigate the KG. On the other hand, we remove the instructional component in the KG-LLM (ablation) knowledge prompt and the CoT reasoning process from the expected output. This approach stands out for its clarity and simplicity, providing a good comprehension of the KG and facilitating improved prediction accuracy.
As illustrated in Figure 2, our experiments utilize the previously described prompts: the KG-LLM knowledge prompt and the KG-LLM (ablation) knowledge prompt. The KG-LLM (ablation) knowledge prompt pairs a question ($Q$) with a target answer ($A$), reflecting the conventional format employed in language models. Conversely, the KG-LLM prompt enriches $Q$ with the KG CoT prompt and embeds the CoT reasoning process within $A$. It further introduces an instruction within $Q$ that narrows down the model’s response options (yes/no), thereby guiding it towards more precise outcomes by constraining the range of possible answers.
In addition, we adopt the approach of utilizing one-shot ICL learning, specifically tailored to the FLAN-T5-Large, which is our smallest model. This is because for models of this scale, the impact of utilizing one-shot ICL versus few-shot ICL on accuracy is minimal [3]. To maintain consistency across our experimental framework, we apply the same one-shot ICL methodology to all three models under investigation: FLAN-T5-Large, LlaMa2-7B, and Gemma-7B. This uniform approach ensures that our comparative analysis of the models’ performances is conducted under equivalent learning conditions, thereby providing a fair and controlled assessment of their capabilities.
@@ 3.3 KG-LLM Framework
Our complete KG-LLM Framework is illustrated in Figure 1. In the initial stage, the original KG is taken as input. The preprocess function randomly selects a unique path in the KG and converts it into the KG-LLM and KG-LLM (ablation) knowledge prompt. During the fine-tuning phase, three distinct LLMs are utilized: Flan-T5-Large, LlaMa2-7B, and Gemma-7B. Different fine-tuning techniques are applied for each model within our framework. A global fine-tuning strategy is employed on Flan-T5 to boost its performance. For LlaMa2 and Gemma, a 4-bit quantized LoRA (Low-Rank Adaptation) modification [8] is implemented. This technique strategically adjusts model parameters with minimal computational overhead, thereby optimizing their efficiency and effectiveness in handling complex prediction tasks. During the training process, we use the cross-entropy loss function $L$. It calculates the difference between the model’s predicted token probabilities and the actual token probabilities in the expected output sequence. In the following equation, $n$ represents the length of the expected output sequence, $x$ stands for the input instruction, and $y_{i}$ denotes the i-th token in the expected output sequence.
$\small L=-\sum_{i=1}^{n}\log P(y_{i}|x,y_{1},y_{2},...,y_{i-1})$ (1)
To evaluate our KG-LLM Frameworks, we train each model twice. As illustrated in Figure 2, the initial training session employs KG-LLM (ablation) knowledge prompt inputs to establish a baseline. Subsequently, we instruction fine-tune the original models using KG-LLM knowledge prompt inputs. The KG-LLM instruction falls under the classification category. By listing all possible options in the instruction, LLMs can more effectively follow and generate a response from the given choices. For instruction details, please refer to Appendix.
Upon the completion of the training phases, we subject each model to two inference task sets, each comprising two sub-tests: non-In-Context Learning (non-ICL) and In-Context Learning (ICL). The primary set of inference tasks is centered around multi-hop link prediction. Conversely, the secondary set probes the models’ generalization ability in multi-hop relation prediction, particularly with previously unseen prompts. Through pre- and post-ICL evaluation within each task set, we aim to evaluate the impact of ICL integration across both the KG-LLM (ablation) and KG-LLM frameworks.
@@ 4 Experiments
$\bullet$ Q2: Does incorporating ICL enhance model performance on multi-hop link prediction tasks?
@@ 4.1 Experimental Setup
We conduct experiments over $two$ real-world datasets, WN18RR and NELL-995, which are constructed and released by the OpenKE library [7]. WN18RR and NELL-995 are two commonly used benchmark datasets for evaluating knowledge graph embedding models in the field of knowledge representation learning. Statistics of the datasets are shown in Table 1.
@@
In the preprocessing stage of the WN18RR and NELL-995 datasets, we generate an equal number of positive and negative instances for training. We proceed by randomly extracting 80% of the entries from both the positive and negative datasets to constitute the training data and 20% for testing. To reduce redundancy in the training data, we eliminate duplicates before combining positive and negative training data. During training, we reserve 20% of the training data as validation data. Balancing the ratio of positive and negative data in prediction tasks is based on the premise that it can enhance the overall performance of the model. This approach acknowledges that real-world knowledge graphs often contain a significantly higher number of negative edges (i.e., absence of relationships) compared to positive edges (i.e., presence of relationships). By ensuring equal representation of positive and negative examples, the model is encouraged to learn to predict both types of edges, preventing it from being biased towards only predicting negative edges and improving its performance in predicting positive edges.
@@
To assess the effectiveness of our KG-LLM frameworks, we compare their performance across all experiments. In the non-ICL multi-hop link prediction testing, we also compare our LLMs approach with traditional approaches such as TransE [2], Analogy [10], CompleX [22], DistMult [33], and RESCAL [13] to evaluate performance.
@@
We trained each model for 2 epochs on an A40 GPU, and despite limited resources, models still showed promising results. In our study, we set the maximum complexity of multi-hop problems to 10 nodes. We also monitor the input token size to optimize processing efficiency, noting that Flan-T5, with its 512-token capacity, had the smallest token size. Consequently, we tailored our experiments to ensure that the maximum length of input data did not exceed 512 tokens.
@@
Datasets WN18RR NELL-995
Metrics F1 AUC F1 AUC
TransE 0.37 0.37 0.26 0.13
Analogy 0.61 0.52 0.29 0.20
CompleX 0.60 0.51 0.29 0.19
DistMult 0.56 0.48 0.25 0.14
Rescal 0.61 0.50 0.59 0.53
Flan-T5-Large (ablation) 0.63 0.67 0.60 0.66
LlaMa2-7B (ablation) 0.74 0.72 0.71 0.73
Gemma-7B (ablation) 0.76 0.73 0.72 0.71
Flan-T5-Large (KG-LLM) 0.73 0.71 0.70 0.72
LlaMa2-7B (KG-LLM) 0.82 0.83 0.81 0.80
Gemma-7B (KG-LLM) 0.84 0.81 0.82 0.83
@@
Datasets WN18RR NELL-995
Metrics F1 AUC F1 AUC
Flan-T5-Large (ablation) 0.70 0.74 0.61 0.63
LlaMa2-7B (ablation) 0.83 0.85 0.81 0.80
Gemma-7B (ablation) 0.86 0.87 0.82 0.83
Flan-T5-Large (KG-LLM) 0.85 0.87 0.78 0.80
LlaMa2-7B (KG-LLM) 0.97 0.95 0.96 0.93
Gemma-7B (KG-LLM) 0.98 0.94 0.95 0.95

### v6

@@ Abstract
The task of multi-hop link prediction within knowledge graphs (KGs) stands as a challenge in the field of knowledge graph analysis, as it requires the model to reason through and understand all intermediate connections before making a prediction. In this paper, we introduce the Knowledge Graph Large Language Model (KG-LLM), a novel framework that leverages large language models (LLMs) for knowledge graph tasks. We first convert structured knowledge graph data into natural language and then use these natural language prompts to fine-tune LLMs to enhance multi-hop link prediction in KGs. By converting the KG to natural language prompts, our framework is designed to learn the latent representations of entities and their interrelations. To show the efficacy of the KG-LLM Framework, we fine-tune three leading LLMs within this framework, including Flan-T5, LLaMa2 and Gemma. Further, we explore the framework’s potential to provide LLMs with zero-shot capabilities for handling previously unseen prompts. Experimental results show that KG-LLM significantly improves the models’ generalization capabilities, leading to more accurate predictions in unfamiliar scenarios. Our code is available at https://anonymous.4open.science/r/KG-LLM-FED0.
@@ 1 Introduction
In the domain of data representation and organization, knowledge graphs (KGs) have emerged as a structured and effective methodology, attracting substantial interest in recent years. Although two-node link prediction in KGs has yielded promising results, the multi-hop link prediction remains a difficult task. In real life, multi-hop link prediction plays a crucial role because, often, we are more interested in the relationship between two far-apart entities rather than direct connections. This requires models to reason through intermediate entities and their relationships. A further challenge is the issue of debugging KG model predictions, particularly in the context of discriminative prediction, where the model’s lack of explanatory reasoning steps, obscures the origins of errors, diminishing accuracy and performance. Consequently, the development of models capable of generatively and precisely predicting multi-hop links in KGs is a critical challenge.
Figure 1: A visual overview of KG-LLM framework
Historically, approaches to solving tasks related to KGs can trace their origins from embedding-based methods to more recent advancements with LLMs [28]. Initially, embedding-based methods played a crucial role, utilizing techniques to represent both entities and relations in a KG as low-dimensional vectors to address the link prediction task by preserving the structural and semantic integrity of the graph [2, 30, 5, 13]. As the field progressed, the integration of LLMs began to offer new paradigms, leveraging large amounts of data and advanced architectures to further enhance prediction capabilities and semantic understanding in KGs [1, 38, 39, 37, 22]. This transition shows a significant improvement from using purely mathematical representations to more context-aware methodologies that better understand the knowledge representation.
@@
To bridge these gaps, we propose the Knowledge Graph Large Language Model (KG-LLM), a novel approach to multi-hop link prediction. As illustrated in Figure 1, nodes in KGs are interconnected through specific relations. Initially, our framework takes input from the original KG dataset. After preprocessing, all paths in the KG will transform into chain-of-thought (CoT) prompts [33], each includes a series of relational statements, which can be represented as {Node 1 (has relation_$x$ with) Node 2, Node 2 (has relation_$y$ with) Node 3, etc.}. The complexity of the multi-hop problem is determined by its path length and the number of nodes. Through instruction fine-tuning (IFT) [31] of three Large Language Models (LLMs): Flan-T5 [31], LlaMa2 [24], and Gemma [23], our framework is ready to enhance multi-hop link prediction performance during the testing phase. Moreover, by integrating in-context learning (ICL) [35], the model not only improves but also has the capacity to tackle unseen prompts, showcasing our method’s innovativeness in addressing multi-hop link prediction challenges.
Our study presents the KG-LLM framework as an innovative approach to the multi-hop link prediction task. Our key contributions are:
@@
  • Our analysis of real-world datasets confirms that our framework improves generative multi-hop link prediction in KGs, underscoring the benefits of incorporating CoT and instruction fine-tuning during training.
@@ 2 Related Work
Recently, researchers have used Graph Neural Network (GNN) models to solve various graph-related tasks, significantly advancing the field. Among different GNN models, Graph Attention Networks (GATs) have gained attention for their ability to weigh the importance of neighboring nodes, with models like wsGAT [6] demonstrating effectiveness in link prediction tasks. Additionally, Graph Convolutional Network (GCN)-based models have shown promising results; ConGLR [12] leverages context graphs and logical reasoning for improved inductive relation prediction, while ConvRot [11] integrates relational rotation and convolutional techniques to enhance link prediction performance in knowledge graphs (KGs). While the aforementioned approaches have achieved significant success, multi-hop link prediction remains an unsolved challenge.
Other than GNN models, recent development of large language models (LLMs), such as BERT [4], GPT [18], LLaMA [24], Gemini [23], and Flan-T5 [31] has also solved various KGs tasks, including link prediction. The text-to-text training approach makes LLMs particularly suitable for our generative multi-hop link prediction task. Recent studies, concurrent work such as GraphEdit [8], MuseGraph [21], and InstructGraph [27], have shown that natural language is effective for representing structural data for LLMs. Besides, training on large-scale data makes it possible for LLMs to generalize to unseen tasks or prompts that were not part of its training data [32].
Another advantage of LLM-based generative modeling is the Chain-of-Thought (CoT) reasoning ability [34]. It provides the flexibility of modifying the instruction, options, and exemplars to allow structured generation and prediction. The Chain-of-Thought reasoning process can be naturally integrated with KGs by translating a reasoning path on a KG into natural language. This flexibility allows us to easily test the model’s ability to follow instructions and make decisions based on the provided information. Similarly, In-Context Learning (ICL) [3] helps LLMs learn from demonstrative examples in the prompt to generate correct answers for the given question. This can also be naturally integrated with KGs. As a result, CoT and ICL enable flexible KG reasoning through natural language.
Figure 2: An Example of Prompt Used in the Multi-hop Link Prediction Training Process: Models processed through the ablation framework will be trained using the ablation knowledge prompt (left), whereas models processed via the KG-LLM framework will be trained on the KG-LLM knowledge prompt (right).
@@ 3.1.1 Multi-hop Link Prediction
The task of multi-hop link prediction extends beyond simple link prediction between two nodes. It aims to identify missing connections over multiple relational steps within a knowledge graph. Specifically, given a sequence of observed triples that form a connected path $P_{obs}={(e_{1},r_{1},e_{2}),(e_{2},r_{2},e_{3}),\ldots,(e_{n-1},r_{n-1},e_{n})}\subseteq L$, where each triple $(e_{i},r_{i},e_{i+1})$ denotes an observed relation $r_{i}$ between entities $e_{i}$ and $e_{i+1}$. The objective is to predict whether a missing link $l_{miss}=(e_{1},?,e_{n})$ exists by answering True or False [19, 26].
@@ 3.1.2 Multi-hop Relation Prediction
The task of multi-hop relation prediction closely aligns with the concept of multi-hop link prediction, with a critical distinction in the question and output. Rather than determining the existence of a missing link $l_{miss}$, this task predicts the relationship. This change shifts the focus from a binary existence query to identifying the specific relationship that binds the entities [16].
@@ 3.2 Knowledge Prompt
The knowledge prompt is a specialized prompt designed for KGs that converts a given sequence of observed triples $P_{obs}$ into natural language. By leveraging the knowledge prompt in the training process, the model can more effectively understand the underlying relationships and patterns present within KGs, thus improving overall performance in multi-hop prediction tasks. In Figure 3, we define the two types of knowledge prompts, KG-LLM knowledge prompt and KG-LLM (ablation) knowledge prompt for both multi-hop link prediction and multi-hop relation prediction.
Figure 3: Overview of our knowledge prompts in the ablation and KG-LLM Frameworks: Ablation framework’s knowledge prompts are in the first and third columns. KG-LLM framework’s knowledge prompts are in the second and fourth columns.
The two types of prompts demonstrate distinct approaches to enhancing model performance in multi-hop prediction tasks. KG-LLM knowledge prompt adopts a structured format that includes instructions and inputs. This approach involves textualizing node and relation IDs into text based on the dataset and breaking down complex inputs into manageable, concise processes. The KG-LLM instruction falls under the classification category. By listing all possible options in the instructions, LLMs can follow and generate a response based on the given choices. On the other hand, we remove the instruction and textualized IDs in the ablation knowledge prompt and the CoT reasoning process from the expected response. This approach stands out for its clarity and simplicity, providing a good comprehension of the KG and improving prediction accuracy. To illustrate our knowledge prompt better, we provide an example for the multi-hop link prediction task in Figure 2.
In addition, we adopt the approach of utilizing one-shot ICL learning, specifically tailored to the FLAN-T5-Large, which is our smallest model. This is because, for models of this scale, the impact of utilizing one-shot ICL versus few-shot ICL on accuracy is minimal [3]. To maintain consistency across our experimental framework, we apply the same one-shot ICL methodology to all LLMs. This uniform approach ensures that our comparative analysis of the models’ performances is conducted under equivalent learning conditions. We listed all ICL examples in Appendix A.1.
@@ 3.3 KG-LLM Framework
Our complete KG-LLM Framework is illustrated in Figure 1. Initially, the KG is taken as input. Each node is iteratively assigned as the root, and depth-first search (DFS) is used to extract all possible paths. Duplicate paths are then removed, retaining only those with node counts ranging from 2 to 6. This range is based on the “six degrees of separation” theory [7], which states that any two individuals are, on average, connected through a chain of no more than six intermediaries. The node counts correspond to the number of hops: a single-hop is between two nodes, a two-hop involves three nodes, and so on. These paths are labeled as either positive (there is a connection between the first and last node) and negative (there is no connection) instances. We observed that negative instances outnumbered positive instances, so we randomly reduced the number of negative instances to achieve a balanced dataset. Finally, these paths are converted into KG-LLM and KG-LLM (ablation) knowledge prompts. During the fine-tuning phase, three distinct LLMs are utilized: Flan-T5-Large, LlaMa2-7B, and Gemma-7B. We added all the node IDs and relation IDs as special tokens to the vocabulary of these LLMs. Different fine-tuning techniques are applied for each model within our framework. A global fine-tuning strategy is employed on Flan-T5 to boost its performance. For LlaMa2 and Gemma, a 4-bit quantized LoRA (Low-Rank Adaptation) modification [10] is implemented. During the training process, we use the cross-entropy loss function $L$. It calculates the difference between the model’s predicted token probabilities and the actual token probabilities in the expected output sequence. In the following equation, $n$ represents the length of the expected output sequence, $x$ stands for the input instruction, and $y_{i}$ denotes the i-th token in the expected output sequence.
$L=-\sum_{i=1}^{n}\log P(y_{i}|x,y_{1},y_{2},...,y_{i-1})$ (1)
To evaluate our KG-LLM Frameworks, we train each model twice. As illustrated in Figure 2, the initial training session employs KG-LLM (ablation) knowledge prompt inputs to establish a baseline. Subsequently, we use instruction finetuning to finetune the original models using KG-LLM knowledge prompt inputs.
After the training phases, we subject each model to two inference task sets, each comprising two sub-tests: non-In-Context Learning (non-ICL) and In-Context Learning (ICL). The primary set of inference tasks is centered around multi-hop link prediction. Conversely, the secondary set probes the models’ generalization ability in multi-hop relation prediction, particularly with previously unseen prompts. Through pre- and post-ICL evaluation within each task set, we aim to evaluate the impact of ICL integration across both the KG-LLM (ablation) and KG-LLM frameworks.
@@ 4 Experiments
$\bullet$ Q2: Does incorporating ICL enhance model performance on multi-hop link prediction task?
@@ 4.1 Experimental Setup
We conduct experiments over $four$ real-world datasets, WN18RR, NELL-995, FB15k-237 and YAGO3-10, which are constructed by the OpenKE library [9]. All datasets are commonly used for evaluating knowledge graph models in the field of knowledge representation learning. Statistics of the datasets are shown in Table 1.
@@
In the preprocessing stage of each dataset, we randomly selected 80% of the nodes to construct the training set of KG. Following the steps in section 3.3, we constructed the training knowledge prompts. For validation, we randomly split off 20% of the positive and negative instances from training knowledge prompts. The same procedure was applied to the remaining 20% of the nodes to create the test set.
@@
To assess the effectiveness of our KG-LLM frameworks, we compare models’ performance across 4 experiments: non-ICL multi-hop link prediction, one-shot ICL multi-hop link prediction, non-ICL multi-hop relation prediction, and one-shot ICL multi-hop relation prediction. In the non-ICL multi-hop link prediction testing, we compare our approach with several traditional methods:
  • TransE [2] is a traditional distance model that represents relationships as translations in the embedding space.
  • Analogy [14] can effectively capture knowledge graph structures to improve link prediction.
  • CompleX [25] uses complex embeddings to represent both entities and relations, capturing asymmetric relationships.
  • DistMult [36] represents relations as diagonal matrices for simplicity and efficiency.
  • RESCAL [17] uses a tensor factorization method that captures rich interactions between entities and relations.
  • wsGAT [6] is a graph attention network that uses weighted self-attention mechanisms to perform various knowledge graph tasks.
  • ConGLR [12] leverages context-aware graph representation learning to enhance link prediction.
  • ConvRot [11] integrates convolutional networks and rotational embeddings to perform a variety of knowledge graph tasks.
@@
We trained each model for 5 epochs on an A40 GPU, and despite limited resources, models still showed promising results. As mentioned in section 3.3, we set the maximum complexity of five-hops. We also monitor the input token size to optimize processing efficiency, noting that Flan-T5, with its 512-token capacity, had the smallest token size. Consequently, we tailored our experiments to ensure that the maximum length of input data did not exceed 512 tokens.
@@
Datasets WN18RR NELL-995 FB15k-237 YAGO3-10
Metrics F1 AUC F1 AUC F1 AUC F1 AUC
TransE 0.37 0.47 0.26 0.43 0.29 0.48 0.34 0.51
Analogy 0.61 0.52 0.29 0.47 0.35 0.54 0.39 0.56
CompleX 0.60 0.51 0.29 0.49 0.32 0.51 0.36 0.53
DistMult 0.56 0.48 0.25 0.44 0.30 0.50 0.35 0.52
Rescal 0.61 0.50 0.59 0.53 0.43 0.58 0.47 0.61
wsGAT 0.69 0.71 0.62 0.67 0.50 0.63 0.54 0.66
ConGLR 0.74 0.69 0.66 0.64 0.63 0.68 0.69 0.61
ConvRot 0.75 0.77 0.72 0.66 0.67 0.62 0.62 0.63
Flan-T5 (Ablation) 0.63 0.67 0.60 0.66 0.63 0.67 0.68 0.70
LLaMa2 (Ablation) 0.74 0.72 0.71 0.73 0.69 0.72 0.76 0.75
Gemma (Ablation) 0.76 0.73 0.72 0.71 0.65 0.73 0.78 0.76
Flan-T5 (KG-LLM) 0.73 0.71 0.70 0.72 0.66 0.70 0.74 0.75
LLaMa2 (KG-LLM) 0.82 0.83 0.81 0.80 0.73 0.76 0.85 0.86
Gemma (KG-LLM) 0.84 0.81 0.82 0.83 0.79 0.81 0.82 0.83
(a) wsGAT
(b) ConGLR
(c) ConvRot
(d) Flan-T5 (Ablation)
(e) LLaMa2 (Ablation)
(f) Gemma (Ablation)
(g) Flan-T5 (KGLLM)
(h) LLaMa2 (KGLLM)
(i) Gemma (KGLLM)
Figure 4: Linear Relationship Between Complexity of Multi-Hop and Performance Score
@@
Datasets WN18RR NELL-995 FB15k-237 YAGO3-10
Metrics F1 AUC F1 AUC F1 AUC F1 AUC
Flan-T5 (Ablation) 0.70 0.74 0.61 0.63 0.55 0.63 0.57 0.63
LLaMa2 (Ablation) 0.83 0.85 0.81 0.80 0.76 0.78 0.84 0.83
Gemma (Ablation) 0.86 0.87 0.82 0.83 0.75 0.80 0.88 0.87
Flan-T5 (KG-LLM) 0.85 0.87 0.68 0.70 0.74 0.77 0.69 0.73
LLaMa2 (KG-LLM) 0.97 0.95 0.96 0.93 0.85 0.86 0.95 0.93
Gemma (KG-LLM) 0.98 0.94 0.95 0.95 0.94 0.92 0.91 0.94